# Localized Distributional Robustness in Submodular Multi-Task Subset Selection

### v2

@@ 2 Preliminaries and Background
While Greedy is demonstrably optimal and very easy to implement, its standard form is usually avoided due to the computational cost of reevaluating every remaining element at each iteration that it incurs. Several methods have been proposed to circumvent this shortcoming, two significant ones being Lazy Greedy[44], and Stochastic Greedy[45]. In particular, Stochastic Greedy achieves a reduced computational cost by reducing its evaluations to a subset of the remaining elements sampled uniformly at random at each iteration, instead of reevaluating every element. Stochastic Greedy also enjoys, on expectation, an approximation ratio similar to that of Greedy, with an additional term that designates the dependence on the cardinality of the randomly sampled subset.
Theorem 1 (Stochastic Greedy approximation ratio).
[45] Let $f\vcentcolon 2^{N}\rightarrow\mathbb{R}$ be a normalized, monotone nondecreasing submodular function. Let $R=(\lvert N\rvert/K)\log(1/\epsilon)$ be the size of the sampled set at each iteration of Stochastic Greedy used in the solution of Problem (1), where $\epsilon\in(0,1)$. Then, on expectation, Stochastic Greedy achieves an approximation ratio of $1-1/e-\epsilon$, i.e., Stochastic Greedy produces a solution $\hat{S}$ which satisfies
$\mathbb{E}[f(\hat{S})]\geq(1-1/e-\epsilon)f(S^{\ast}),$ (9)
where $S^{\ast}$ is a maximizer of Problem (1).
In the formulation of Problem (6), we have intentionally left unspecified the statistical distance $D(P\;\|\;Q)$ to be used in the regularization term, in order to localize our region of interest to a neighborhood around the reference distribution $Q$. In this section, we will be specifically interested in investigating which choices of $D(\cdot\;\|\;\cdot)$ lead to favorable algorithmic solutions and theoretical guarantees for Problem (6). In particular, of specific interest will be such $D(\cdot\;\|\;\cdot)$ that preserve the submodular nature of the problem. For, ensuring that the problem remains submodular allows us to produce a locally robust solution essentially at no additional cost with respect to the non-robust formulation of Problem (1), by falling back to our familiar tools such as Stochastic Greedy, and enjoying theoretical guarantees such as that presented in Theorem 1.
We first shift our attention to metrics induced by $\ell^{p}$ norms. One choice for $D(P\;\|\;Q)$ that leads to a favorable algorithmic solution is the metric induced by the $\ell^{\infty}$ norm, i.e., $D(P\;\|\;Q)=\lVert P-Q\rVert_{\infty}\vcentcolon=\max_{i\in[n]}\lvert P_{i}-Q_{i}\rvert.$ With this selection, Problem (6) becomes
$\begin{split}&\max_{S\subseteq N}\min_{P\in\Delta_{n}}\sum^{n}_{i=1}P_{i}f^{i}(S)+\lambda\lVert P-Q\rVert_{\infty}\\ &\text{s. t. }\lvert S\rvert\leq K.\end{split}$ (10)
In this case, we may exploit the reduction of the inner problem of (10) to a linear program to derive a result. Indeed, since we know that the optimal point of a linear program will occur at one of the vertices of the feasible region [46], we have the following equivalence:
$\begin{split}\min_{P\in\Delta_{n}}\sum^{n}_{i=1}P_{i}f^{i}(S)+\lambda\lVert P-Q\rVert_{\infty}\equiv\min_{i\in[n]}f^{i}(S)-\lambda Q_{i}.\end{split}$ (11)
Hence, Problem (10) reduces, in whole, to
$\begin{split}&\max_{S\subseteq N}\min_{i\in[n]}f^{i}(S)-\lambda Q_{i}\\ &\text{s. t. }\lvert S\rvert\leq K.\end{split}$ (12)
The main observation about Problem (12) is that it is strikingly similar to the worst-case formulation of Problem (2), however, in addition, the information due to the reference distribution $Q$ still holds influence. We propose to use SSA [19] with a slight modification enabling the incorporation of the reference distribution $Q$ for the solution of Problem (12). The full algorithm, which we name Saturate with Preference, is presented in Algorithm 1. It has a function evaluation complexity of $\mathcal{O}(\lvert N\rvert^{2}n\log(n\min_{i\in[n]}f^{i}(N))$, identical to the standard SSA [19].
It is worth noting that the choice of the metric induced by the $\ell^{1}$ norm, i.e., $D(P\;\|\;Q)=\lVert P-Q\rVert_{1}\vcentcolon=\sum_{i=1}^{n}\lvert P_{i}-Q_{i}\rvert$, through the same reduction of the inner problem to a linear program, and the same equivalence of (11), results in the same formulation of Problem (12). Since $P$ and $Q$ are discrete probability distributions, one may also view the quantity $\lVert P-Q\rVert_{1}$ as a scaled version of the total variation distance, an example of an flimit-$f-$divergence, which motivates the discussion of the next section, in using one flimit-$f-$divergence with particularly significant theoretical outcomes, namely, the relative entropy or KL-divergence.
Algorithm 1 Saturate with Preference
Input: Finite collection of monotone nondecreasing submodular functions $f^{1},\ldots f^{n}$, ground set $N$, cardinality constraint $K\in\mathbb{N}$, regularization parameter $\lambda>0$, reference distribution $Q\in\Delta_{n}$
   Output: Solution set $S$
1:$k_{m}\leftarrow 0$
2:$k_{M}\leftarrow\min_{i}f^{i}(N)-\lambda Q_{i}$
3:$S\leftarrow\emptyset$
@@ 3 Investigation of Various Statistical Distances for Regularization
5:   $k\leftarrow(k_{M}-k_{m})/2$
6:   Define $\bar{f}_{(k)}(\mathord{\cdot})\leftarrow\frac{1}{n}\sum_{i=1}^{n}\min\{f^{i}(\mathord{\cdot})-\lambda Q_{i},k\}$
7:   $\hat{S}\leftarrow\textsc{Greedy}(\bar{f}_{(k)},k)$
8:   if $\lvert\hat{S}\rvert>\alpha K$ then
@@
4 Relative Entropy Regularization
@@
leads to the most significant theoretical outcomes, as will be seen in the following analysis. With this choice, the novel formulation is as follows:
$\begin{split}&\max_{S}\min_{P}\sum^{n}_{i=1}P_{i}f^{i}(S)+\lambda D_{\text{KL}}(P\;\|\;Q)\\ &\text{s. t. }S\subseteq N,|S|\leq K\text{, and }P\in\Delta_{n}.\end{split}$ (14)
We start our analysis of the novel formulation by considering the dual formulation of the inner minimization problem in (14), by introducing a Lagrange multiplier for the constraint $P\in\Delta_{n}$ in Problem (14). We first note that this constraint is equivalent to the two constraints (i) $P_{i}\geq 0$ for all $i\in[n]$ and (ii) $\sum_{i=1}^{n}P_{i}=1.$ Introducing a Lagrange multiplier only for constraint (ii), we write the Lagrangian of the inner problem as
$\mathcal{L}\left(P,S,\mu\right)=\sum_{i=1}^{n}P_{i}f^{i}(S)+\lambda D_{\text{KL}}(P\;\|\;Q)+\mu\sum_{i=1}^{n}P_{i}-\mu.$ (15)
Minimizing $\mathcal{L}(P,S,\mu)$ over $P$ will yield the dual formulation of the inner problem of (14). We solve
$\begin{split}&\min_{P}\mathcal{L}\left(P,S,\mu\right)\\ &\text{s. t. }P_{i}\geq 0\text{ for all }i\in[n].\end{split}$ (16)
@@
$\frac{\partial\mathcal{L}}{\partial P_{i}}=f^{i}(S)+\lambda\log\left(\frac{P_{i}}{Q_{i}}\right)+\lambda+\mu.$ (17)
Setting this quantity equal to 00, we obtain the minimizer $P_{i}^{*}$:
$P_{i}^{*}=Q_{i}\exp{\left(-\frac{\lambda+\mu+f^{i}(S)}{\lambda}\right)}.$ (18)
@@
$\begin{split}\sum_{i=1}^{n}P_{i}^{*}&=\sum_{i=1}^{n}Q_{i}\exp{\left(-\frac{\lambda+\mu+f^{i}(S)}{\lambda}\right)}\\ &=\sum_{i=1}^{n}Q_{i}\exp{\left(-\frac{f^{i}(S)}{\lambda}\right)}\exp{\left(-\frac{\lambda+\mu}{\lambda}\right)}=1.\end{split}$ (19)
@@
$\lambda+\mu=\lambda\log\left(\sum_{i=1}^{n}Q_{i}\exp{\left(\frac{-f^{i}(S)}{\lambda}\right)}\right).$ (20)
@@
$\begin{split}\mathcal{L}(P^{*},S,\mu)&=\sum_{i=1}^{n}P_{i}^{*}f^{i}(S)+\lambda D_{\text{KL}}(P^{*}\;\|\;Q)\\ &=\sum_{i=1}^{n}Q_{i}\exp{\left(-\frac{\lambda+\mu+f^{i}(S)}{\lambda}\right)}f^{i}(S)\\ &+\lambda\sum_{i=1}^{n}Q_{i}\exp{\left(-\frac{\lambda+\mu+f^{i}(S)}{\lambda}\right)}\cdot\left[-\frac{\lambda+\mu+f^{i}(S)}{\lambda}\right]\\ &=-\left(\lambda+\mu\right)\sum_{i=1}^{n}Q_{i}\exp{\left(-\frac{\lambda+\mu+f^{i}(S)}{\lambda}\right)}.\\ \end{split}$ (21)
@@
$\begin{split}\sum_{i=1}^{n}Q_{i}\exp{\left(-\frac{\lambda+\mu+f^{i}(S)}{\lambda}\right)}=\sum_{i=1}^{n}P_{i}^{*}=1,\end{split}$ (22)
@@
$\begin{split}\mathcal{L}(P^{*},S,\mu)=-\left(\lambda+\mu\right).\end{split}$ (23)
@@
$\begin{split}G(S)\vcentcolon=\mathcal{L}(P^{*},S,\mu)=-\lambda\log\left(\sum_{i=1}^{n}Q_{i}\exp{\left(\frac{-f^{i}(S)}{\lambda}\right)}\right).\end{split}$ (24)
Hence, the dual formulation of Problem (14) becomes that of the maximization of another set function $G$:
$\begin{split}&\max_{S\subseteq N}G(S)\\ &\text{s.t. }\lvert S\rvert\leq K.\end{split}$ (25)
A simple analysis of the extreme cases of values for the regularization parameter $\lambda$ is revealing of the equivalence of the two problems. We evaluate the limit
$\begin{split}\lim_{\lambda\to x}-\lambda\log\left(\sum_{i=1}^{n}Q_{i}\exp{\left(\frac{-f^{i}(S)}{\lambda}\right)}\right)=\lim_{\lambda\to x}-\frac{\log\left(\sum_{i=1}^{n}Q_{i}\exp{\left(\frac{-f^{i}(S)}{\lambda}\right)}\right)}{\frac{1}{\lambda}}\end{split}$ (26)
@@
$\lim_{\lambda\to x}\frac{\sum_{i=1}^{n}Q_{i}\exp{\left(\frac{-f^{i}(S)}{\lambda}\right)f^{i}(S)}}{\sum_{i=1}^{n}Q_{i}\exp{\left(\frac{-f^{i}(S)}{\lambda}\right)}}.$ (27)
Letting $x\to\infty$, we have:
$\begin{split}\lim_{\lambda\to\infty}\frac{\sum_{i=1}^{n}Q_{i}\exp{\left(\frac{-f^{i}(S)}{\lambda}\right)f^{i}(S)}}{\sum_{i=1}^{n}Q_{i}\exp{\left(\frac{-f^{i}(S)}{\lambda}\right)}}=\sum_{i=1}^{n}Q_{i}f^{i}(S).\end{split}$ (28)
This is indeed the expected behavior, since letting $\lambda\to\infty$ in Problem (14) effectively assigns all importance to the regularization term, and necessarily forces one to have $P^{\ast}=Q$, making the objective
$\begin{split}&\max_{S\subseteq N}\sum_{i=1}^{n}Q_{i}f^{i}(S)\\ &\text{s.t. }\lvert S\rvert\leq K,\end{split}$ (29)
@@
$\begin{split}&P_{i}^{\ast}=Q_{i}\cdot\exp\left(-\dfrac{\lambda\log\left(\sum_{j=1}^{n}Q_{j}\exp\left(\dfrac{-f^{j}(S)}{\lambda}\right)\right)+f^{i}(S)}{\lambda}\right).\end{split}$ (30)
This means that for any finite value of $\lambda$, for $P^{\ast}=Q$, we need, for all $i\in[n]$,
$\exp\left(-\dfrac{\lambda\log\left(\sum_{j=1}^{n}Q_{j}\exp\left(\dfrac{-f^{j}(S)}{\lambda}\right)\right)+f^{i}(S)}{\lambda}\right)=1,$ (31)
which, in turn, means,
$\lambda\log\left(\sum_{j=1}^{n}Q_{j}\exp\left(\dfrac{f^{j}(S)}{\lambda}\right)\right)=-f^{i}(S),\forall i\in[n],$ (32)
and this last condition implies that $f^{i}(S)=f^{j}(S)$, for any two $i,j\in[n]$, i.e., for any finite value of $\lambda$, the values of all objectives must have the same value at $S$ for $P^{\ast}=Q$. Outside of this very specific circumstance, at no finite value of $\lambda$ does $P^{\ast}=Q$ hold.
On the other hand, letting $x=0$, the value of the limit is determined by the most dominant term in the sums in the numerator and the denominator, that is,
$\begin{split}&\lim_{\lambda\to 0}\frac{\sum_{i=1}^{n}Q_{i}\exp{\left(\frac{-f^{i}(S)}{\lambda}\right)f^{i}(S)}}{\sum_{i=1}^{n}Q_{i}\exp{\left(\frac{-f^{i}(S)}{\lambda}\right)}}=\min_{i\in[n]}\frac{Q_{i}\exp{\left(\frac{-f^{i}(S)}{\lambda}\right)f^{i}(S)}}{Q_{i}\exp{\left(\frac{-f^{i}(S)}{\lambda}\right)}}=\min_{i\in[n]}f^{i}(S).\end{split}$ (33)
@@
$\begin{split}&\max_{S\subseteq N}\min_{i\in[n]}f^{i}(S)\\ &\text{s.t. }\lvert S\rvert\leq K.\end{split}$ (34)
Again, letting $x=0$ in Problem (14) removes the regularization term, and reduces the inner minimization problem to a linear program, whose solution will have $P_{i}=1$ for $i=\arg\!\min_{i\in[n]}f^{i}(S)$, and $P_{j}=0$ for all $j\neq i$, exactly recovering (34). Note that this is also exactly the worst-case formulation of Problem (2), demonstrating that our novel formulation is a generalization of the worst-case formulation.
The main question that arises directly from this novel formulation of Problem (25) is whether the set function $G$ retains the properties of being normalized, monotone nondecreasing, and submodular since the presence of all three of these properties would reduce the solution of Problem (25) to the use of standardized methods, such as Stochastic Greedy. Fortunately, we are not far from this ideal case, as demonstrated in Theorem 2.
@@
Input: Monotone nondecreasing submodular function $f$, ground set $N$, cardinality constraint $K\in\mathbb{N}$, random sampling set size $r$
   Output: Solution set $S$
1:$S\leftarrow\emptyset$
2:while $|S|<K$ do
3:   Sample a subset $R$ of size $r$ uniformly at random from $N\setminus S$
4:   $S\leftarrow S\cup\arg\!\max_{e\in R}\Delta_{f}(e|S)$
@@
The set function $G(S)$ may be expressed as the composition of two functions, $G(S)=g(h(S))$, where the set function
$h(S)\vcentcolon=\sum_{i=1}^{n}Q_{i}\left(1-\exp\left(-\frac{f^{i}(S)}{\lambda}\right)\right)$ (35)
is
  1. (i)
    normalized,
  2. (ii)
    monotone nondecreasing, and
  3. (iii)
    submodular,
and the function $g(x)\vcentcolon=-\lambda\log(1-x)$ is
  1. (iv)
    monotone increasing,
  2. (v)
    convex, and
  3. (vi)
    Lipschitz continuous.
in $\operatorname{dom}g=[0,h(S^{\ast})]$.
@@
    $\begin{split}h(\emptyset)&=\sum_{i=1}^{n}Q_{i}\left(1-\exp{\left(-\frac{f^{i}(\emptyset)}{\lambda}\right)}\right)\\ &=\sum_{i=1}^{n}Q_{i}\left(1-\exp{\left(0\right)}\right)\\ &=0.\end{split}$ (36)
@@
    Let $S\subseteq T\subseteq N$. Then, for each $i\in[n],$ we have $f^{i}(S)\leq f^{i}(T).$ Hence, $-f^{i}(S)/\lambda\geq-f^{i}(T)/\lambda,$ and $\exp(-f^{i}(S)/\lambda)\geq\exp(-f^{i}(T)/\lambda),$ thanks to the monotone increasing property of the $\exp$ function and the positivity of $\lambda$. This, in turn, means that
    $1-\exp\left(-\frac{f^{i}(S)}{\lambda}\right)\leq 1-\exp\left(-\frac{f^{i}(T)}{\lambda}\right),$ (37)
@@
    $\begin{split}\sum_{i=1}^{n}Q_{i}\left(1-\exp\left(-\frac{f^{i}(S)}{\lambda}\right)\right)\leq\sum_{i=1}^{n}Q_{i}\left(1-\exp\left(-\frac{f^{i}(T)}{\lambda}\right)\right),\end{split}$ (38)
    that is, $h(S)\leq h(T)$.
@@
    Let $S\subseteq T\subset N$, $e\in N\setminus T$. For all $i\in[n]$, we have:
    $\dfrac{f^{i}(S\cup\{e\})-f^{i}(S)}{\lambda}\geq\dfrac{f^{i}(T\cup\{e\})-f^{i}(T)}{\lambda}.$ (39)
@@
    $\begin{split}1-\exp\left(\dfrac{-f^{i}(S\cup\{e\})+f^{i}(S)}{\lambda}\right)\geq 1-\exp\left(\dfrac{-f^{i}(T\cup\{e\})+f^{i}(T)}{\lambda}\right).\end{split}$ (40)
    Furthermore, because each $f^{i}$ is monotone nondecreasing, $f^{i}(S)\leq f^{i}(T)$, and since $x\mapsto\exp(-x/\lambda)$ is a monotone decreasing function for $\lambda>0$, we have
    $\exp\left(\frac{-f^{i}(S)}{\lambda}\right)\geq\exp\left(\frac{-f^{i}(T)}{\lambda}\right).$ (41)
    Furthermore, both of these quantities are nonnegative. Then, multiplying the left-hand side of (40) by $\exp(-f^{i}(S)/\lambda)$ and the right-hand side by $\exp(-f^{i}(T)/\lambda)$ preserves the inequality, and we obtain
    $\begin{split}\exp\left(\frac{-f^{i}(S)}{\lambda}\right)-\exp\left(\frac{-f^{i}(S\cup\{e\})}{\lambda}\right)\\ \geq\exp\left(\frac{-f^{i}(T)}{\lambda}\right)-\exp\left(\frac{-f^{i}(T\cup\{e\})}{\lambda}\right).\end{split}$ (42)
    This demonstrates that the set function $S\mapsto-\exp\left(-f^{i}(S)/\lambda\right)$ is submodular. The addition of the constant $1$, multiplication with the nonnegative $Q_{i}$ and the summation through the indices $i\in[n]$ preserve the submodularity, hence,
    $h(S)=\sum_{i=1}^{n}Q_{i}\left(1-\exp\left(\dfrac{-f^{i}(S)}{\lambda}\right)\right)$ (43)
@@
    for all $\lambda>0$ and $x\in[0,h(S^{\ast})]$. Furthermore, it is of note that $\frac{\mathrm{d}g}{\mathrm{d}x}$ attains its maximum value at $x=h(S^{\ast})$.

### v3

@@ 3 Investigation of Various Statistical Distances for Regularization
[…]
@@
[…]
6: Define $\bar{f}_{(k)}(\mathcal{X}):=\frac{1}{n}\sum_{i=1}^{n}\min\{f_{i}(\mathcal{X})-\lambda Q_{i},\;k\}$
[…]
8:if $\lvert\hat{\mathcal{S}}\rvert>\alpha K$ then
@@
15:return $\mathcal{S}$
4 Relative-Entropy Regularization
@@
leads to the most significant theoretical outcomes, as we will see in the following analysis. With this choice, the formulation becomes:
$\begin{gathered}\max_{\mathcal{S}\subseteq\mathcal{N}}\min_{P\in\Delta_{n}}\sum^{n}_{i=1}P_{i}f_{i}(\mathcal{S})+\lambda D_{\text{KL}}(P\;\|\;Q)\\ \text{s.t.}|\mathcal{S}|\leq K.\end{gathered}$ (14)
We consider the dual of the inner minimization problem, by introducing a Lagrange multiplier for the constraint $P\in\Delta_{n}$. This constraint is equivalent to the two constraints (i) $P_{i}\geq 0$ for all $i\in[n]$ and (ii) $\sum_{i=1}^{n}P_{i}=1.$ Introducing a Lagrange multiplier only for constraint (ii), we write the Lagrangian of the inner problem as
$\mathcal{L}\left(P,\mathcal{S},\mu\right)=\sum_{i=1}^{n}P_{i}f_{i}(\mathcal{S})+\lambda D_{\text{KL}}(P\;\|\;Q)+\mu\sum_{i=1}^{n}P_{i}-\mu.$ (15)
Minimizing $\mathcal{L}(P,\mathcal{S},\mu)$ over $P$ will yield the dual formulation of the inner problem of (14). We solve
$\begin{gathered}\min_{P\in\mathbb{R}^{n}}\mathcal{L}\left(P,\mathcal{S},\mu\right)\\ \text{s.t.}\;P_{i}\geq 0\text{ for all }i\in[n].\end{gathered}$ (16)
@@
$\frac{\partial\mathcal{L}}{\partial P_{i}}=f_{i}(\mathcal{S})+\lambda\log\left(\frac{P_{i}}{Q_{i}}\right)+\lambda+\mu.$ (17)
Setting this equal to 0, we obtain the minimizer $P_{i}^{*}$:
$P_{i}^{*}=Q_{i}\exp{\left(-\frac{\lambda+\mu+f_{i}(\mathcal{S})}{\lambda}\right)}.$ (18)
@@
$\begin{split}\sum_{i=1}^{n}P_{i}^{*}&=\sum_{i=1}^{n}Q_{i}\exp{\left(-\frac{\lambda+\mu+f_{i}(\mathcal{S})}{\lambda}\right)}=\sum_{i=1}^{n}Q_{i}\exp{\left(-\frac{f_{i}(\mathcal{S})}{\lambda}\right)}\exp{\left(-\frac{\lambda+\mu}{\lambda}\right)}=1.\end{split}$ (19)
@@
$\lambda+\mu=\lambda\log\left(\sum_{i=1}^{n}Q_{i}\exp{\left(\frac{-f_{i}(\mathcal{S})}{\lambda}\right)}\right).$ (20)
@@
[…]
@@
$\begin{split}\sum_{i=1}^{n}Q_{i}\exp{\left(-\frac{\lambda+\mu+f_{i}(\mathcal{S})}{\lambda}\right)}=\sum_{i=1}^{n}P_{i}^{*}=1,\end{split}$ (22)
@@
$\begin{split}\mathcal{L}(P^{*},\mathcal{S},\mu)=-\left(\lambda+\mu\right).\end{split}$ (23)
@@
$\begin{split}G(\mathcal{S})\vcentcolon=\mathcal{L}(P^{*},\mathcal{S},\mu)=-\lambda\log\left(\sum_{i=1}^{n}Q_{i}\exp{\left(\frac{-f_{i}(\mathcal{S})}{\lambda}\right)}\right).\end{split}$ (24)
Hence, the dual formulation of (14) becomes that of the maximization of another set function $G$:
$\begin{gathered}\max_{\mathcal{S}\subseteq\mathcal{N}}G(\mathcal{S})\\ \text{s.t.}\;\lvert\mathcal{S}\rvert\leq K.\end{gathered}$ (25)
A simple analysis of the extreme cases of $\lambda$ provides good intuition on the equivalence of the two problems. Let us evaluate the limit
$\begin{split}\lim_{\lambda\to x}-\lambda\log\left(\sum_{i=1}^{n}Q_{i}\exp{\left(\frac{-f_{i}(\mathcal{S})}{\lambda}\right)}\right)=\lim_{\lambda\to x}-\frac{\log\left(\sum_{i=1}^{n}Q_{i}\exp{\left(\frac{-f_{i}(\mathcal{S})}{\lambda}\right)}\right)}{\frac{1}{\lambda}}.\end{split}$ (26)
@@
$\lim_{\lambda\to x}\frac{\sum_{i=1}^{n}Q_{i}\exp{\left(\frac{-f_{i}(\mathcal{S})}{\lambda}\right)f_{i}(\mathcal{S})}}{\sum_{i=1}^{n}Q_{i}\exp{\left(\frac{-f_{i}(\mathcal{S})}{\lambda}\right)}}.$ (27)
Letting $x\to\infty$, we have
$\begin{split}\lim_{\lambda\to\infty}\frac{\sum_{i=1}^{n}Q_{i}\exp{\left(\frac{-f_{i}(\mathcal{S})}{\lambda}\right)f_{i}(\mathcal{S})}}{\sum_{i=1}^{n}Q_{i}\exp{\left(\frac{-f_{i}(\mathcal{S})}{\lambda}\right)}}=\sum_{i=1}^{n}Q_{i}f_{i}(\mathcal{S}).\end{split}$ (28)
This is indeed the expected behavior, since letting $\lambda\to\infty$ in (14) effectively assigns all importance to the regularization term and forces one to have $P^{\ast}=Q$, making the objective
$\begin{gathered}\max_{\mathcal{S}\subseteq\mathcal{N}}\sum_{i=1}^{n}Q_{i}f_{i}(\mathcal{S})\\ \text{s.t.}\;\lvert\mathcal{S}\rvert\leq K,\end{gathered}$ (29)
@@
$\begin{split}&P_{i}^{\ast}=Q_{i}\cdot\exp\left(-\dfrac{\lambda\log\left(\sum_{j=1}^{n}Q_{j}\exp\left(\dfrac{-f_{j}(\mathcal{S})}{\lambda}\right)\right)+f_{i}(\mathcal{S})}{\lambda}\right).\end{split}$ (30)
This means that for any finite value of $\lambda$, to get $P^{\ast}=Q$, we need, for all $i\in[n]$,
$\exp\left(-\dfrac{\lambda\log\left(\sum_{j=1}^{n}Q_{j}\exp\left(\dfrac{-f_{j}(\mathcal{S})}{\lambda}\right)\right)+f_{i}(\mathcal{S})}{\lambda}\right)=1,$ (31)
which, in turn, means
$\lambda\log\left(\sum_{j=1}^{n}Q_{j}\exp\left(\dfrac{f_{j}(\mathcal{S})}{\lambda}\right)\right)=-f_{i}(\mathcal{S})$ (32)
for all $i\in[n]$. This last condition implies that $f_{i}(\mathcal{S})=f_{j}(\mathcal{S})$, for any two $i,j\in[n]$, i.e., for any finite value of $\lambda$, the values of all objectives must have the same value at $\mathcal{S}$ for $P^{\ast}=Q$. Outside of this very specific circumstance, at no finite value of $\lambda$ does $P^{\ast}=Q$ hold.
On the other hand, letting $x=0$, the value of the limit is determined by the most dominant terms in the sums in the numerator and the denominator, that is,
$\begin{split}&\lim_{\lambda\to 0}\frac{\sum_{i=1}^{n}Q_{i}\exp{\left(\frac{-f_{i}(\mathcal{S})}{\lambda}\right)f_{i}(\mathcal{S})}}{\sum_{i=1}^{n}Q_{i}\exp{\left(\frac{-f_{i}(\mathcal{S})}{\lambda}\right)}}=\min_{i\in[n]}\frac{Q_{i}\exp{\left(\frac{-f_{i}(\mathcal{S})}{\lambda}\right)f_{i}(\mathcal{S})}}{Q_{i}\exp{\left(\frac{-f_{i}(\mathcal{S})}{\lambda}\right)}}=\min_{i\in[n]}f_{i}(\mathcal{S}).\end{split}$ (33)
@@
$\begin{gathered}\max_{\mathcal{S}\subseteq\mathcal{N}}\min_{i\in[n]}f_{i}(\mathcal{S})\\ \text{s.t.}\;\lvert\mathcal{S}\rvert\leq K.\end{gathered}$ (34)
Again, letting $x=0$ in (14) removes the regularization term, and reduces the inner minimization problem to a linear program, whose solution will have $P_{i}=1$ for $i=\arg\!\min_{i\in[n]}f_{i}(\mathcal{S})$, and $P_{j}=0$ for all $j\neq i$, producing (34). Note that this is also exactly the worst-case formulation of (2), demonstrating that our formulation is a generalization of the worst-case formulation.
The main question that arises directly from the formulation of (25) is whether the set function $G$ retains the properties of being normalized, monotone nondecreasing, and submodular since the presence of all three of these properties would reduce its solution to the use of standardized methods, such as Stochastic Greedy. Fortunately, we are not far from this ideal case, as demonstrated in Theorem 2.
@@
Input: Monotone nondecreasing submodular function $f$, ground set $\mathcal{N}$, integer cardinality bound $0\leq K\leq\lvert\mathcal{N}\rvert$, random sampling set size $r$
Output: Solution set $\mathcal{S}$
[…]
2:while $|\mathcal{S}|<K$ do
3: Sample a subset $\mathcal{R}$ of size $r$ uniformly at random from $\mathcal{N}\setminus\mathcal{S}$
[…]
@@
The set function $G(\mathcal{S})$ can be expressed as the composition of two functions, $G(S)=g(h(\mathcal{S}))$, where the set function
$h(\mathcal{S})\vcentcolon=\sum_{i=1}^{n}Q_{i}\left(1-\exp\left(-\frac{f_{i}(\mathcal{S})}{\lambda}\right)\right)$ (35)
is (i) normalized, (ii) monotone nondecreasing, and (iii) submodular, and the function $g(x)\vcentcolon=-\lambda\log(1-x)$ is (iv) monotone increasing, (v) convex, and (vi) Lipschitz continuous in $\operatorname{dom}g=[0,h(\mathcal{S}^{\ast})]$.
@@
    $\begin{split}h(\emptyset)=\sum_{i=1}^{n}Q_{i}\left(1-\exp{\left(-\frac{f_{i}(\emptyset)}{\lambda}\right)}\right)=\sum_{i=1}^{n}Q_{i}\left(1-\exp{\left(0\right)}\right)=0.\end{split}$ (36)
@@
    Let $\mathcal{S}\subseteq\mathcal{T}\subseteq\mathcal{N}$. Then, for each $i\in[n],$ we have $f_{i}(\mathcal{S})\leq f_{i}(\mathcal{T}).$ Hence, $f_{i}(\mathcal{S})/\lambda\geq-f_{i}(\mathcal{T})/\lambda,$ and $\exp(-f_{i}(\mathcal{S})/\lambda)\geq\exp(-f_{i}(\mathcal{T})/\lambda),$ thanks to the monotone increasing property of the $\exp$ function and the positivity of $\lambda$. This means that
    $1-\exp\left(-\frac{f_{i}(\mathcal{S})}{\lambda}\right)\leq 1-\exp\left(-\frac{f_{i}(\mathcal{T})}{\lambda}\right),$ (37)
@@
    $\begin{split}\sum_{i=1}^{n}Q_{i}\left(1-\exp\left(-\frac{f_{i}(\mathcal{S})}{\lambda}\right)\right)\leq\sum_{i=1}^{n}Q_{i}\left(1-\exp\left(-\frac{f_{i}(\mathcal{T})}{\lambda}\right)\right),\end{split}$ (38)
    that is, $h(\mathcal{S})\leq h(\mathcal{T})$.
@@
    Let $\mathcal{S}\subseteq\mathcal{T}\subset\mathcal{N}$, $e\in\mathcal{N}\setminus\mathcal{T}$. For all $i\in[n]$, we have
    $\dfrac{f_{i}(\mathcal{S}\cup\{e\})-f_{i}(\mathcal{S})}{\lambda}\geq\dfrac{f_{i}(\mathcal{T}\cup\{e\})-f_{i}(\mathcal{T})}{\lambda}.$ (39)
@@
    $\begin{split}1-\exp\left(\dfrac{-f_{i}(\mathcal{S}\cup\{e\})+f_{i}(\mathcal{S})}{\lambda}\right)\geq 1-\exp\left(\dfrac{-f_{i}(\mathcal{T}\cup\{e\})+f_{i}(\mathcal{T})}{\lambda}\right).\end{split}$ (40)
    Furthermore, because each $f_{i}$ is monotone nondecreasing, $f_{i}(\mathcal{S})\leq f_{i}(\mathcal{T})$, and since $x\mapsto\exp(-x/\lambda)$ is a monotone decreasing function for $\lambda>0$, we have
    $\exp\left(\frac{-f_{i}(\mathcal{S})}{\lambda}\right)\geq\exp\left(\frac{-f_{i}(\mathcal{T})}{\lambda}\right).$ (41)
    Both of these quantities are nonnegative. Then, multiplying the left-hand side of (40) by $\exp(-f_{i}(\mathcal{S})/\lambda)$ and the right-hand side by $\exp(-f_{i}(\mathcal{T})/\lambda)$ preserves the inequality, and we obtain
    $\begin{split}\exp\left(\frac{-f_{i}(\mathcal{S})}{\lambda}\right)-\exp\left(\frac{-f_{i}(\mathcal{S}\cup\{e\})}{\lambda}\right)\geq\exp\left(\frac{-f_{i}(\mathcal{T})}{\lambda}\right)-\exp\left(\frac{-f_{i}(\mathcal{T}\cup\{e\})}{\lambda}\right).\end{split}$ (42)
    This demonstrates that the set function $\mathcal{S}\mapsto-\exp(-f_{i}(\mathcal{S})/\lambda)$ is submodular for all $i\in[n]$. The addition of the constant $1$, multiplication with the nonnegative $Q_{i}$ and the summation through the indices $i\in[n]$ preserve submodularity, hence,
    $h(\mathcal{S})=\sum_{i=1}^{n}Q_{i}\left(1-\exp\left(\dfrac{-f_{i}(\mathcal{S})}{\lambda}\right)\right)$ (43)
@@
    for all $\lambda>0$ and $x\in[0,h(\mathcal{S}^{\ast})]$.